# Bridging Quantum Computing and Nuclear Structure: Atomic Nuclei on a Trapped-Ion Quantum Computer

### v1

@@ Methods
Figure 4: Quantum circuits for pUCCD ansatz. Sketched quantum circuits for preparing the pUCCD ansatz taking 20O as an example. a. Circuit layout suitable for devices with heavy-hexagon connectivity. b. Circuit layout for all-to-all connectivity. First, the reference state is prepared by applying $X$ gates to qubits. Then, Givens rotations (Eq. (16)) are applied to take into account particle-hole excitations of nucleon pairs. Single-qubit rotations are applied as needed prior to measuring Hamiltonian expectation values.
A related study by Sarma et al. [16] considered oxygen isotopes in the $sd$ shell on a trapped-ion quantum computer, IonQ Aria, using a variant of unitary coupled cluster doubles (UCCD) ansatz applied only to time-reversal nucleon pairs (denoted UCCD ($\nu=0$) in the following where $\nu$ is the seniority).
Although formally equivalent to DOCI, their implementation yielded ground-state errors of $3-13\%$ relative to FCI, depending on the isotope. One limitation arises from the scaling of circuit parameters: the number of Givens rotations required grows rapidly with system size, making direct application to heavier isotopes impractical.
$N_{\mathrm{param}}$ $=N_{\mathrm{occ.}}(N_{q}-N_{\mathrm{occ.}}),$ (11)
@@
where the last expression corresponds to a first-order Trotter approximation with a single step, which we adopt in this work.
Each factor in Eq. (15) is equivalent to a Givens rotation acting on qubits $p$ and $h$:
$G(\theta)$ $=\begin{pmatrix}1&0&0&0\\ 0&\cos(\theta/2)&-\sin(\theta/2)&0\\ 0&\sin(\theta/2)&\cos(\theta/2)&0\\ 0&0&0&1\end{pmatrix}$ (16)
Thus, the pUCCD ansatz can be implemented using relatively shallow circuits.
Fig. 4 illustrates two circuit layouts for preparing the pUCCD ansatz, taking 20O nucleus as an example. a. pUCCD (GS): optimized for devices with limited connectivity (e.g., IBM heavy-hexagon architectures), using Givens rotations followed by SWAP gates (GS) [25]. The basic idea behind the circuit is detailed in Ref. [50]. b. pUCCD (G): one realization for all-to-all connectivity (e.g., trapped-ion devices), requiring only Givens rotations.
@@
While pUCCD (GS) has a nearly unique circuit layout, pUCCD (G) admits many realizations depending on the choice of initial configuration and ordering of Givens rotations. We therefore present simulator results for pUCCD (G) as distributions over multiple circuit realizations. In what follows, we use a single deterministic realization for both pUCCD (GS) and pUCCD (G) unless otherwise noted. Adaptive strategies such as orbital optimization or ansatz refinement could further enhance accuracy, but are left for future work.
Implementation of Givens rotation gates
To construct the pUCCD ansatz circuits (Fig. 4) and to perform basis rotations for diagonalizing the $XX+YY$ terms, it is necessary to implement Givens rotation gates.
Several approaches are known in literature. In this work, we examined the following three methods:
@@
At the raw circuit level, these methods require 2, 4, and 3 CNOT gates per Givens rotation, respectively. However, we found the number of CNOTs for the transpiled circuits are equivalent to each other. As a result, the total numbers of CNOTs in the for pUCCD (GS) and pUCCD (G) ansatze scale as $3N_{G}$ and $2N_{G}$, where $N_{G}$ is the number of Givens rotations in the ansatz.
@@
Quantum circuits submitted to the hardware are transformed to native gate sets of the adopted hardware. For the RIKEN-Quantinuum’s Reimei device, CNOT gates are to be transpiled to another two-qubit gate, the so-called ZZPhase gate, $\exp(-i\pi\alpha/2(Z_{p}\otimes Z_{q}))$. Here we follow the notation of pytket [53, 54], i.e., angle parameter is given in the unit of $\pi/2$. Indeed, the number of ZZPhase gates for the compiled pUCCD (G) ansatz became twice the number of Givens rotations listed in Tab. 3.
Measurement of energy expectation values
In this work, the parameters of the pUCCD ansatze were optimized on statevector simulators using the Adam optimizer using PennyLane [55]. This is still an idealized setting, but good starting point to evaluate the ansatz quality and measurement strategies. The strategy to optimize the ansatz circuit on a real device is left as a future issue. Optimization directly on quantum hardware remains an open problem, for which gradient-free sequential approaches [56] may offer a promising path.
Diagonal terms in the computational basis, i.e., operators involving $\{I,Z,ZZ\}$, can be obtained by measuring the ansatz circuit in the computational basis. In this case, post-selection techniques is naturally applicable: bitstrings violating the particle number of the target system are discarded.
In contrast, non-diagonal terms $\{XX,YY\}$ require additional circuits. This is especially costly on hardware with limited connectivity, where extra SWAP gates increase circuit depth and accumulate noise. We therefore considered two measurement strategies: First one, referred to as the “Hadamard" method in Results section, is the straightforward approach using Hadamard gates. The ansatz is followed by Hadamard gates on each qubit, after which measurements are taken in the computational basis. This allows direct access to all $X_{p}X_{q}$ expectation values. Since $XX$ and $YY$ terms contribute equally to the energy, $\langle YY\rangle$ are evaluated by doubling $\langle XX\rangle$ values. While straightforward, this method does not preserve particle-number symmetry, limiting the use of post-selection. Note that one can consider applying the so-called computational basis sampling method [57]. The second strategy is to be referred to as basis rotation (BasisRot), diagonalizing $XX+YY$ terms prior to measurement [29, 25]:
$\left\langle{\psi}\right|$ $X_{p}X_{q}+Y_{p}Y_{q}\left|{\psi}\right\rangle$
where $\mathcal{U}_{p,q}(\theta)$ is a Givens rotation $G(2\theta)$, Eq. (16), applied to qubits $p$ and $q$. Following Ref. [29], the required terms can be grouped such that only $2\lfloor N_{q}/2\rfloor$ additional circuits are needed to measure all the $XX+YY$ terms, where $N_{q}$ is the number of qubits. On devices with all-to-all connectivity, SWAP gates for this arrangement are not strictly necessary; however, we employed them to make comparison with the H-gate method rather straightforward. Importantly, this strategy preserves particle-number symmetry, enabling effective post-selection.
It should be noted that energies estimated from separately measured terms is not guaranteed to be lower than the true ground state energy.are generally non-variational, even with post-selection. Since diagonal and off-diagonal contributions are measured independently, their sum is not guaranteed to remain below the true ground-state energy.
Remarks on statistical uncertainty of hardware results
Figure 5: Statistical uncertainty of hardware results for $^{66}\mathrm{Ni}$. a diagonal terms. b XX+$Y$ terms evaluated by two measurement strategies: Hadamard (green) and basis rotation (red with hatching). The distributions are obtained by bootstrap, i.e., resampling the raw data 500 times. The vertical dashed line shows the ideal statevector simulation result.
To quantify statistical uncertainties in hardware results, we applied a bootstrapping analysis. As an illustration, Fig. 5, shows the distribution of the ground-state energy of $^{66}\mathrm{Ni}$ obtained by resampling the data 500 times. The energy is separated into contributions from diagonal terms (top panel) and $XX+YY$ terms (bottom panel). Vertical dashed lines indicate the corresponding noise-free statevector results for comparison.
The accuracy on the diagonal term depends primarily on occupation numbers, while the $XX+YY$ terms are contribution is more sensitive to both occupation numbers and relative phases between configurations. As shown in the lower panel of Fig. 5, the distribution obtained using the Hadamard method is noticeably broader than that obtained with basis rotations. This indicates that the basis-rotation strategy is more robust against statistical noise, especially when the number of measurement shots is limited.
Hardware information
@@
The experiments on the Reimei trapped-ion quantum computer were carried out via the Quantinuum Nexus [28] platform. Table 4 summarizes the calibration data of the Reimei device, which were obtained on 2025-05-09. The reported metrics include single- and two-qubit gate errors, SPAM (state preparation and measurement) errors, crosstalk, and memory error rates. These parameters provide a baseline for evaluating the quality of quantum circuits executed on the device. In particular, the relatively low one- and two-qubit gate error rates, combined with the all-to-all connectivity of trapped-ion hardware, enable the accurate implementation of the pUCCD ansatz circuits described in this work.
@@
Table 4: Calibration data of the Reimei device The values are from data on 2025-05-09 via the Quantinuum’s Nexus [28] platform.
@@
Data Availability
The data presented in this manuscript are available from the GitHub repository [58].
Code availability

### v2

@@ II.1 Target systems and Hamiltonian mapping
Figure 2: Sketched quantum circuits for preparing the pUCCD ansatz taking 22O as an example. Original single-particle states in the $sd$ shell consist of twelve states for neutrons, which are now represented by six qubits in the HCB representation. (a) Circuit layout suitable for devices with heavy-hexagon connectivity. (b) Circuit layout for all-to-all connectivity. First, the reference state is prepared by applying $X$ gates to qubits. Then, Givens rotations are applied to take into account particle-hole excitations of nucleon pairs. Single-qubit rotations are applied as needed prior to measuring Hamiltonian expectation values.
The paired-neutron operators satisfy bosonic commutation relations, as detailed in e.g. [33, 21, 20], but multiple occupancy of the same pair is forbidden, leading to the term hard-core boson. A key advantage of the HCB representation—beyond the qubit reduction—is the elimination of non-locality. From Jordan-Wigner mappings, Pauli-$Z$ terms appear to maintain the anti-commutation relations of the fermionic operators. In some literature, these Pauli-$Z$ terms are omitted on purpose, which is referred to as qubit excitation in Refs. [36, 37], as a hardware-efficient mapping. However, in the HCB representation, those Pauli-$Z$ terms naturally cancel out. Hence, one can map the Hamiltonian to the Pauli operators into simple forms as
$H_{\mathrm{HCB;q}}$ $=\sum_{i}\frac{\varepsilon_{i}+V_{ii}}{2}(I_{i}-Z_{i})+\frac{1}{4}\sum_{i\neq j}V_{ij}(X_{i}X_{j}+Y_{i}Y_{j})$
$+\frac{1}{4}\sum_{i\neq j}V^{m}_{ij}(I_{i}-Z_{i})(I_{j}-Z_{j}),$ (10)
All terms of the Hamiltonian Eq. (10) except the second are diagonal in the computational basis and can be measured directly by measuring the expectation values of the Pauli-$Z$ operators. One needs additional circuits to measure the non-diagonal terms $XX+YY$, which will be discussed later in Sec. II.3.
II.2 ansatze
The ansatz, circuit for state preparation, plays a central role in quantum algorithms both in the NISQ and FTQC regimes. In practice, the ansatz must satisfy two requirements: (1) it should be implementable on available quantum hardware, and (2) it should be expressive enough to approximate, e.g., the true ground state with high fidelity.
In this work, we employ the pair unitary coupled cluster doubles (pUCCD) ansatz for even-even multi-neutron systems within the HCB representation. The pUCCD ansatz is a quantum version of the pair coupled cluster doubles method. A seminal work on the pUCCD ansatz in the context of quantum computing is found in Ref. [21], where strategies for measurement and post-selection were also proposed. In chemistry, pUCCD has been successfully applied to systems of paired electrons with spin up and spin down [21, 22, 23]. For nuclear systems, the pUCCD ansatz has not been explored previously (to our knowledge), and our work represents the first application of this approach to nuclear structure calculations.
@@
where the last expression corresponds to a first-order Trotter approximation with a single step, which we adopt in this work. Each terms in Eq. (14) can be implemented as a Givens rotation gate acting on qubits $p$ and $h$:
$G(2\theta)$ $=\begin{pmatrix}1&0&0&0\\ 0&\cos{\theta}&-\sin{\theta}&0\\ 0&\sin{\theta}&\cos{\theta}&0\\ 0&0&0&1\end{pmatrix}$. (15)
Thus, the pUCCD ansatz can be implemented as a sequence of Givens rotations on top of a reference state, e.g., initially prepared by applying $X$ gates to the qubits corresponding to occupied pair orbitals. Fig. 2 illustrates two circuit layouts for preparing the pUCCD ansatz, taking 22O nucleus as an example. (a) pUCCD(GS): optimized for devices with limited connectivity (e.g., IBM heavy-hexagon architectures), using Givens rotations followed by SWAP gates (GS) [21]. The basic idea behind the circuit is detailed in Ref. [38]. (b) pUCCD(G): one realization for all-to-all connectivity (e.g., trapped-ion devices) using only Givens rotations (G). In both cases, the number of parameters, angles of Givens rotations, is the same, given by the number of particle-hole pairs.
However, the circuit depth and number of two-qubit gates slightly differ. While pUCCD(GS) has a nearly unique circuit layout, pUCCD(G) admits many realizations depending on the choice of initial configuration and ordering of Givens rotations. We will therefore present simulator results for pUCCD(G) as distributions over multiple circuit realizations in Sec. III. In what follows, we use a single deterministic realization for both pUCCD(GS) and pUCCD(G) unless otherwise noted.
As a related study, it should be noted that a work by Sarma et al. [15] considered oxygen isotopes in the $sd$ shell on a trapped-ion quantum computer, IonQ Aria, using a variant of unitary coupled cluster doubles (UCCD) ansatz. Their UCCD ansatz was applied only to time-reversal nucleon pairs (denoted UCCD($\nu=0$) in the following where $\nu$ is the seniority). Although formally equivalent to DOCI, their implementation yielded ground-state errors of $3-13\%$ relative to FCI, depending on the isotope. One limitation arises from the scaling of circuit parameters: the number of Givens rotations to express the excitations/de-excitations of nucleon pairs grows rapidly with system size, making direct application to heavier isotopes impractical.
$N_{\mathrm{param}}$ $=N_{\mathrm{occ.}}(N_{q}-N_{\mathrm{occ.}}),$ (16)
@@
II.3 Measurement of energy expectation values
Once the ansatz circuit is prepared, the next task is to evaluate the expectation value of the Hamiltonian. First, we discuss our optimization approach; then we detail the measurement strategies. In this work, the parameters of the pUCCD ansatze were optimized on statevector simulators using the Adam optimizer with PennyLane [39]. This is still an idealized setting, but good starting point to evaluate the ansatz quality and measurement strategies, and to benchmark hardware results. The strategy to optimize the ansatz circuit on a real device is left as a future issue. Optimization directly on quantum hardware remains an open problem, for which gradient-free sequential approaches [40] may offer a promising path.
Diagonal terms in the computational basis, i.e., operators involving $\{I,Z_{i},Z_{i}\otimes Z_{j}\}$, can be obtained by measuring the ansatz circuit in the computational basis. In this case, post-selection techniques is naturally applicable: bitstrings violating the particle number of the target system are discarded.
In contrast, non-diagonal terms $\{X_{i}\otimes X_{j},Y_{i}\otimes Y_{j}\}$ require additional circuits. Both $XX$ and $YY$ terms are local, i.e., free from $Z$-strings, and contribute equally to the energy expectation value. Thus, all $X_{p}X_{q}$ terms can be evaluated by a single additional circuit, ansatz followed by Hadamard gates on all qubits prior to measurement, and $\langle Y_{p}Y_{q}\rangle$ terms can be evaluated by doubling the $\langle X_{p}X_{q}\rangle$ results. This approach will be referred to as the Hadamard method in the following. While straightforward, this method does not preserve particle-number symmetry, limiting the use of post-selection. The other approach employed in this work is one to be referred to as Basis rotations, which involves diagonalizing $XX+YY$ terms in the computational basis prior to measurement [41, 21]:
$\left\langle{\psi}\right|$ $X_{p}X_{q}+Y_{p}Y_{q}\left|{\psi}\right\rangle$
where $\mathcal{U}_{p,q}(\theta)$ can be implemented as a Givens rotation $G(2\theta)$ on qubits $p$ and $q$. In accordance with the discussion in Ref. [41], the required terms can be grouped such that only $2\lfloor N_{q}/2\rfloor$ additional circuits are needed to measure all the $XX+YY$ terms, where $N_{q}$ is the number of qubits. Importantly, this strategy preserves particle-number symmetry, enabling effective post-selection.
Note that since we measure diagonal and off-diagonal contributions in separate circuits, the summed estimate of the energy is not strictly variational even after post-selection. However, in practice, we observed that the energy estimates distributed around the noise-free values from statevector simulations, hence non-variational nature does not pose significant issues in this work.
II.4 Implementation of Givens rotation gates
To construct the pUCCD ansatz circuits (Fig. 2) and to perform basis rotations for diagonalizing the $XX+YY$ terms, it is necessary to implement Givens rotation gates. Several approaches are known in literature and we examined the following three methods:
@@ II.4 Implementation of Givens rotation gates
At the raw circuit level, these methods require 2, 4, and 3 CNOT gates per Givens rotation, respectively. However, we found the number of corresponding 2-qubit native gates for the transpiled/compiled circuits are equivalent to each other. As a result, the total number of the 2-qubit gates for pUCCD(GS) and pUCCD(G) ansatze scales as $3N_{G}$ and $2N_{G}$, respectively. Here, $N_{G}$ is the number of Givens rotations. Thus, the choice of the implementation method can be made based on convenience and the first method was adopted in this work.
@@
Quantum circuits submitted to the hardware are transformed to native gate sets of the adopted hardware. On the RIKEN-Quantinuum’s Reimei device, two-qubit native gates are ZZPhase gates, $\exp(-i\pi\alpha/2(Z_{p}\otimes Z_{q}))$. Here we follow the notation of pytket [44, *pytket-quantinuum], i.e., angle parameter is given in the unit of $\pi/2$. In practice, the compiled circuit uses exactly twice as many ZZPhase gates as there are Givens rotations.
II.5 Hardware information
@@
We now describe the quantum hardware used in this work. The Reimei is based on Quantinuum’s H-1 series trapped-ion quantum computer, and has 20 qubits with all-to-all connectivity. The Reimei device was installed in RIKEN’s Wako campus, and started its operation in February 2025. Table 3 summarizes the calibration data of the Reimei device, which were measured on 2025-05-09. The reported metrics include single- and two-qubit gate errors, SPAM (state preparation and measurement) errors, crosstalk, and memory error rates. These parameters provide a baseline for evaluating the quality of quantum circuits executed on the device. In particular, the relatively low one- and two-qubit gate error rates, combined with the all-to-all connectivity of trapped-ion hardware, enable the accurate implementation of the pUCCD ansatz circuits described in this work.
The experiments on the Reimei trapped-ion quantum computer of this work were carried out via the Quantinuum Nexus [46] platform, which enables users to run quantum simulations on various backends, including noise-free statevector simulators, emulators, and real devices.
@@
Table 3: Calibration data of the Reimei device. The values are from data on 2025-05-09 via the Quantinuum’s Nexus [46] platform.
@@
III Results
Figure 3: Ground-state energies of target isotopes on noise-free statevector simulators. (a)–(c) Ground-state energies of oxygen, calcium, and nickel isotopes, respectively. Full configuration interaction (FCI, dotted lines with gray circles) and doubly occupied configuration interaction (DOCI, blue diamonds) are compared with statevector simulations using the pUCCD(GS) ansatz (green stars) and the pUCCD(G) ansatz (red triangles). Here, pUCCD(GS) denotes the pUCCD ansatz implemented with nearest-neighbor Givens rotations and SWAP gates, while pUCCD(G) assumes all-to-all connectivity, as realized in the Reimei device. (d)–(f) Differences between the pUCCD ansatz results and the DOCI results are shown relative to FCI in logarithmic scale. Uncertainty bands for pUCCD(G) are obtained by random sampling over circuits with different initial configurations and Givens-rotation orderings.
In this section, we present the results of the pUCCD ansatz both on noise-free statevector simulators and on the Reimei quantum hardware for the ground-state energies of oxygen, calcium, and nickel isotopes. The former allows us to assess the intrinsic accuracy of the pUCCD ansatz, while the latter demonstrates the feasibility of executing the pUCCD ansatz on real quantum hardware.
III.1 Ideal simulation results of the pUCCD ansatz
We first assess the performance of the pUCCD ansatz on a noiseless statevector simulator, thereby probing its intrinsic accuracy in approximating ground-state energies of the target isotopes. Full Configuration Interaction (FCI) provides the exact ground-state energies within the valence space, serving as a benchmark for evaluating the accuracy of approximate methods. Doubly Occupied Configuration Interaction (DOCI) is another approximate method that restricts the configuration space to zero-seniority configurations, i.e., states in which all nucleons are paired as time-reversed partners. In noise-free simulations, any deviation from FCI obtained with the pUCCD ansatz reflects intrinsic limitations of the ansatz itself, independent of hardware noise or statistical noise due to the finite number of measurement shots.
Table 4: Summary of pUCCD results on the RIKEN-Quantinuum Reimei device. FCI and ideal pUCCD are classical results. Reimei columns report hardware energies using Hadamard or basis-rotation measurements; errors are relative to ideal pUCCD. Valid ratio is the fraction of bitstrings passing particle-number post-selection.
FCI Ideal pUCCD Reimei: Hadamard Reimei: Basis rotation
Nuclide Energy (MeV) Energy (MeV) Energy (MeV) Rel. Error (%) Energy (MeV) Rel. Error (%) Valid ratio (%)
18O 11.932-11.932 11.932-11.932 11.802-11.802 $1.09$ 11.893-11.893 $0.33$ $98$
20O 23.632-23.632 22.939-22.939 22.383-22.383 $2.42$ 22.854-22.854 $0.37$ $98$
22O 34.498-34.498 33.924-33.924 33.560-33.560 $1.07$ 33.829-33.829 $0.28$ $97$
24O 41.225-41.225 41.021-41.021 40.985-40.985 $0.09$ 41.027-41.027 $0.02$ $97$
26O 40.869-40.869 40.869-40.869 40.613-40.613 $0.63$ 40.940-40.940 $0.18$ $96$
42Ca 19.734-19.734 19.734-19.734 19.727-19.727 $0.04$ 19.775-19.775 $0.21$ $97$
44Ca 38.675-38.675 38.462-38.462 38.540-38.540 $0.20$ 38.424-38.424 $0.10$ $97$
46Ca 56.667-56.667 56.440-56.440 56.414-56.414 $0.05$ 56.565-56.565 $0.22$ $96$
48Ca 73.662-73.662 73.556-73.556 73.538-73.538 $0.02$ 73.489-73.489 $0.09$ $95$
50Ca 85.055-85.055 84.921-84.921 85.135-85.135 $0.25$ 85.007-85.007 $0.10$ $94$
52Ca 95.360-95.360 95.190-95.190 95.038-95.038 $0.16$ 95.149-95.149 $0.04$ $94$
54Ca 102.632-102.632 102.557-102.557 102.622-102.622 $0.06$ 102.648-102.648 $0.09$ $94$
56Ca 104.589-104.589 104.462-104.462 104.620-104.620 $0.15$ 104.325-104.325 $0.13$ $94$
58Ca 106.666-106.666 106.666-106.666 106.885-106.885 $0.20$ 106.668-106.668 <0.01<0.01 $94$
58Ni 21.447-21.447 21.447-21.447 21.207-21.207 $1.12$ 21.291-21.291 $0.73$ $93$
60Ni 41.276-41.276 40.961-40.961 40.752-40.752 $0.51$ 40.992-40.992 $0.07$ $94$
62Ni 59.602-59.602 59.013-59.013 58.703-58.703 $0.52$ 59.015-59.015 <0.01<0.01 $93$
64Ni 76.444-76.444 75.687-75.687 76.027-76.027 $0.45$ 75.765-75.765 $0.10$ $92$
66Ni 91.941-91.941 91.059-91.059 90.552-90.552 $0.56$ 90.695-90.695 $0.40$ $93$
68Ni 106.140-106.140 105.237-105.237 105.169-105.169 $0.06$ 105.288-105.288 $0.05$ $92$
70Ni 119.034-119.034 118.215-118.215 118.189-118.189 $0.02$ 118.158-118.158 $0.05$ $91$
72Ni 130.930-130.930 130.379-130.379 130.511-130.511 $0.10$ 130.397-130.397 $0.01$ $95$
74Ni 141.932-141.932 141.694-141.694 141.565-141.565 $0.09$ 141.715-141.715 $0.02$ $93$
76Ni 152.103-152.103 152.103-152.103 151.986-151.986 $0.08$ 152.054-152.054 $0.03$ $94$
In what follows, we consider two types of pUCCD ansatz, named pUCCD(GS) and pUCCD(G). The (GS) represents Givens plus SWAP, which consists of only nearest neighbor Givens rotations and thereby more suitable for hardware with a limited connectivity [41, 21]. On the other hand, (G) stands for Givens rotations where all-to-all connectivity is assumed. In experiments on the Reimei device, we will focus on the pUCCD(G) ansatz.
Fig. 3 shows ground-state energies for the target isotopes, obtained with pUCCD(G) and pUCCD(GS) on noise-free statevector simulators compared to FCI and DOCI results. The upper panels (a)–(c) display the ground-state energies. At this scale, all curves are nearly indistinguishable. The lower panels (d)–(f) show the differences between pUCCD and DOCI results relative to FCI in logarithmic scale. Results for two-neutron (or two-neutron-hole) systems are omitted in the lower panels, since both DOCI and pUCCD exactly reproduce FCI in those cases.
For the pUCCD ansatze, the symbols show the results of one representative realization of the ansatz, and the uncertainty bands are plotted for the pUCCD(G) ansatz. We generated many pUCCD(G) realizations by varying the initial configuration and the order of Givens rotations in the ansatz circuit. The resulting energies from a skewed distribution around the representative realization, and the uncertainty bands are constructed by fitting Gamma distributions to the histograms of sampled energies. More technical details are provided in the Supplemental Material [35].
Most importantly, pUCCD reproduces DOCI energies with an error of mostly less than $1\%$, while using far fewer parameters, demonstrating its compact yet accurate description. For reference, DOCI itself reproduces FCI ground-state energies within $0.1-1\%$, with deviations smaller than the typical uncertainties (a few percent) of modern nuclear many-body methods [47]. The residual differences depend on the target system: Ca isotopes in the $pf$ shell are well described within the zero-seniority space, whereas oxygen isotopes show larger contributions from higher-seniority configurations leading to slightly larger deviations from FCI. Nickel isotopes in the $jj45$ shell exhibit larger discrepancies between DOCI and pUCCD compared to oxygen and calcium isotopes. This arises from the factors intrinsic to the $jj45$ shell. The valence space, $jj45$ shell, consists of four $jj$-coupling orbitals ($1p_{1/2},1p_{3/2},0f_{5/2},0g_{9/2}$), and the configuration mixing among these orbitals is pronounced in Ni isotopes even for ground states since their effective single particle energies are close to each other [30]. The single-Trotter pUCCD ansatz, while efficient for capturing pairing correlations, becomes less expressive in the presence of such strong configuration mixing.
III.2 Results on RIKEN-Quantinuum’s Reimei trapped-ion quantum device
Figure 4: Relative errors in ground-state energies on the RIKEN-Quantinuum Reimei device. The results are shown for oxygen (a), calcium (b), and nickel (c) isotopes, using two measurement strategies: Hadamard and basis rotation. The error bars represent 1 $\sigma$ statistical uncertainties evaluated by bootstrapping, and the shaded band corresponds to a $0.3\%$ error.
Figure 5: Bootstrapping of hardware results for $^{66}\mathrm{Ni}$. (a) diagonal terms. (b) $XX+YY$ terms evaluated by two measurement strategies: Hadamard (green) and basis rotation (red with hatching). The vertical dashed line shows the ideal statevector simulation result. The bin width is automatically determined by the Freedman-Diaconis rule.
We now report the hardware results for ground-state energies of oxygen, calcium, and nickel isotopes using the pUCCD(G) ansatz on Reimei. Table 4 summarizes exact FCI energies, the ideal pUCCD results from noise-free statevector simulations, and the experimentally measured energies with two measurement strategies along with percent errors relative to the ideal pUCCD results and valid ratios. The valid ratio indicates the ratio of bitstrings that preserve the correct particle number of the target nucleus after measurement, reflecting the hardware fidelity in preparing the desired state. The numbers are from ansatz measurements, hence independent of the measurement strategy for $XX+YY$ terms.
As a whole, the hardware results agree closely with the noise-free simulations, showing relative errors mostly below $0.3\%$ to thedeal pUCCD results, and at most a few percent relative to FCI. These results demonstrate the capability of current quantum hardware to accurately capture nuclear pairing correlations.
This trend is further supported by bootstrapping analyses. Fig. 4 displays the relative errors of ground-state energies obtained on Reimei for oxygen, calcium, and nickel isotopes, using both measurement strategies. Error bars represent 1 $\sigma$ statistical uncertainties evaluated by bootstrapping with 500 resamplings of datasets obtained from 1,024 measurement shots. The shaded band corresponds to a $0.3\%$ error for reference. Most data points fall within this range.
Note again that the difference between Hadamard and Basis rotation methods lies only in the measurement of $XX+YY$ terms, while diagonal terms are measured in the same way for both methods. Between the two measurement strategies, the basis-rotation method consistently outperforms the Hadamard method, yielding results closer to ideal simulations. Lighter isotopes exhibit larger uncertainties because $XX+YY$ terms contribute to the total energy relatively larger in lighter isotopes. In contrast, for heavier isotopes, the mass-dependent scaling commonly applied to two-body matrix elements in phenomenological shell-model interactions reduces the absolute size of the $XX+YY$ contributions (e.g., TBMEs for 58Ca and 78Ni are multiplied by $\sim 0.9$ compared with 42Ca and 58Ni), which further diminishes their relative contribution to the total energy.
Next, we further examine the statistical uncertainties by separating the total energy into diagonal and $XX+YY$ contributions. As an illustration, Fig. 5, shows the distribution of the ground-state energy of $^{66}\mathrm{Ni}$ obtained by resampling the data 500 times. The energy is separated into contributions from diagonal terms (top panel) and $XX+YY$ terms (bottom panel). Vertical dashed lines indicate the corresponding noise-free statevector results for comparison.
The accuracy on the diagonal term depends primarily on occupation numbers, while the $XX+YY$ terms are contribution is more sensitive to both occupation numbers and relative phases between configurations. As shown in the lower panel of Fig. 5, the distribution obtained using the Hadamard method is noticeably broader than that obtained with basis rotations.
IV Summary and outlook
Quantum many-body systems have become a central focus of quantum computing research over the past decade, driven by rapid progress in algorithms and hardware. However, demonstrating genuine quantum advantage in interacting many-body physics remains an open challenge, requiring advances in all components of the algorithmic workflow: ansatz design, quantum algorithms, measurement strategies, optimization techniques, and so on. This work concentrated on the ansatz and measurement aspects, from the standpoint of nuclear structure, where exceptionally strong correlations make the choice of ansatz especially critical.
By combining a hard-core boson mapping of the nuclear shell-model Hamiltonian with a pair-unitary coupled cluster doubles (pUCCD) ansatz, we demonstrated accurate ground-state calculations of oxygen, calcium, and nickel isotopes on the RIKEN-Quantinuum Reimei trapped-ion quantum computer. Our results achieve sub-percent agreement with noise-free statevector simulations, setting a new benchmark for nuclear-structure simulations on real NISQ hardware.
Although the pUCCD parameters were optimized using classical simulators, this proof-of-principle study highlights how physically motivated, compact ansatze can deliver accurate results with modest qubit numbers and circuit depths. Such strategies may guide the discovery of efficient variational forms for larger systems where classical pre-optimization is infeasible.
Looking ahead, incorporating proton-neutron interactions will require ansatze capable of describing more general non-local correlations beyond pairing. Nonetheless, the present results demonstrate the viability of pUCCD for multi-neutron (or multi-proton) systems, and point toward applications to heavier systems such as neutron drops or neutron-rich isotopes. As an illustration of possible extensions, entropy-driven entanglement forging [48] has recently been proposed to localize correlations in Adapt-VQE circuits; similar ideas could be combined with pUCCD, for instance by monitoring generalized seniority. In addition, hybrid qubit encoding schemes [49] could be extended to nuclear systems.
Beyond ansatz design, new algorithmic paradigms merit exploration. Quantum subspace diagonalization techniques [50, 51, 52, 53, 54], including Quantum Krylov and related methods, use quantum devices to generate subspace states while delegating diagonalization to classical post-processing. These approaches provide access to excited states and transition properties such as electromagnetic transitions. Such hybrid methods, in combination with physically motivated ansatze and developing efficient schemes for time evolution and measurement, may offer a promising path toward realistic quantum simulations of nuclear systems beyond the reach of classical methods.
Data Availability Statement